OpenStudy (anonymous):

If h(x) = −3x − 10, find h(−3).

OpenStudy (anonymous):

h(-3) = -3(-3) - 10 =9-10 =-1

OpenStudy (anonymous):

Thnx bro :)

OpenStudy (anonymous):

Np